OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)?

OpenStudy (bradely):

y-y1 = m(x-x1) y+8= 6(x+1) y+8=6x+6 6x-y-2=0 sourece: http://www.mathskey.com/question2answer/

OpenStudy (nurali):

Start from point-slope form: y-y1=m(x-x1) y+8 = 6(x+1) Simplify: y+8 = 6x+6 Subtract 6x from both sides: -6x + y + 8 = 6 Subtract 8 to both sides: -6x + y = -2 Multiply both sides by -1 to get a positive coefficient on x: 6x - y = 2

OpenStudy (anonymous):

thankyou both of you! I am awful at math.

OpenStudy (nurali):

Anytime.